# IsoBench: Benchmarking Multimodal Foundation Models on Isomorphic Representations

### v2

@@ 1 Introduction
We introduce IsoBench, a test dataset consisting of 1,630 samples222The IsoBench dataset currently consists of 1,630 examples across four domains, though the construction of a larger dataset is currently underway. spanning diverse domains such as discrete and applied mathematics, physics, chemistry, and chess. For each sample, we evaluate multiple isomorphic input representations containing the same information—one with a visual representation and others with domain-specific textual representations—to facilitate multimodal performance assessments.
@@ 4 IsoCombination and IsoScratchPad
Based on the observations made above on IsoBench, we can design two simple mechanisms to further study the performance of multimodal foundation models on tasks with isomorphic representations: IsoCombination (IsoCB) and IsoScratchPad (IsoSP). IsoCB feeds multiple isomorphic representations to the multimodal model simultaneously in order to see the combined effect of using multiple representations on performance. Alternatively, IsoSP uses a scratchpad to explicitly translate from visual to text representations (and then uses this text representation to solve a task), taking advantage of the empirically higher-performing text representations for visual data. In particular, for IsoSP we employ a two-step prompting strategy: initially, the model receives a prompt featuring a visual representation, which it is tasked to translate into a textual format; then, the model is prompted with this generated text representation to predict the output. We illustrate both methods in Figure 3.
@@ 5 Related Works
There has been a plethora of recent developments in vision-language foundation models; they can be broadly categorized by their methods for representing visual modalities. A representative approach involves tokenizing visual inputs to be jointly trained with language inputs (Aghajanyan et al., 2022; Yu et al., 2023; Google, 2023; McKinzie et al., 2024; Chen et al., 2022; 2023, inter alia). Another line of work processes continuous visual features by directly projecting them to the language embedding space via a learnable function (Liu et al., 2023b; a; 2024a; Bavishi et al., 2023). While the design choices of some API-access models (OpenAI, 2023b; Anthropic, 2024; Google, 2023; Reka, 2024) remain largely unknown, most performant models use early fusion, the practice of integrating visual and language features at the input level of an autoregressive model. At the core of these design choices is the hardness in representing visual features, which has been reported by several early studies (McKinzie et al., 2024) to be the key bottleneck towards performant vision-language foundation models. IsoBench aims to supplement prior works, which have mostly focused on analyzing modeling methods, with a data-centric view by measuring performance discrepancies incurred by multimodal representations of the same problem.
@@
Many vision-language datasets have recently been curated to assess the capabilities of multimodal foundation models (Yue et al., 2023; Lu et al., 2024; Zhang et al., 2024a; Fan et al., 2024; Zhang et al., 2024b). They often consist of a large collection of problems with varying levels of difficulty and usage of the visual inputs. As a result, they are suitable for assessing VLMs as a whole, but cannot attribute these performances to capabilities of their LLM backbones and improvements from downstream vision-language training. While several contemporaneous works (Fan et al., 2024; Zhang et al., 2024a; b) introduce a language-only subset of their vision-language prompts, they focus primarily on assessing certain areas of foundation models, such as deductive reasoning and problems of beyond-polynomial complexity. IsoBench is the first of its kind that offers a holistic and fine-grained analysis of performance gap induced by variations in input modalities.
@@
Similar to what IsoBench observes on multimodal foundation models’ sensitivity to input modalities, uni-modal language models also show sensitivity to input perturbation. For example, LLMs are shown to be sensitive to subtle changes in zero-shot and few-show settings (Sclar et al., 2023; Chang & Jia, 2023). Meanwhile, the sensitivity of Transformers, the backbone of foundation models, is also broadly studied (Bombari & Mondelli, 2024; Bhattamishra et al., 2023; Vasudeva et al., 2024). To address the sensitivity issues, methods including prompt design (Yoo et al., 2021; Mishra et al., 2021; Le et al., 2023; Liu et al., 2023c) and prompting-based training strategies (Jain et al., 2023; Guo et al., 2023) have been employed. IsoBench is one of the pioneers to study model sensitivity to input representations and modalities.
@@
Recent work has shown that Transformers, the backbone of multimodal foundation models, are capable of implementing various algorithms after fine-tuning on particular tasks (Khalighinejad et al., 2023; Garg et al., 2022; von Oswald et al., 2022; Hanna et al., 2023; Fu et al., 2023; Lee et al., 2023) or with proper prompting schemes (Wei et al., 2023; Yao et al., 2023; Liu et al., 2024b). However, the extent to which foundation models are capable of solving problems involving complex reasoning, algorithms, or scientific awareness remains unclear.

### v3

@@ 1 Introduction
We introduce IsoBench, a test dataset consisting of 1,887 samples222The IsoBench dataset (www.huggingface.co/datasets/isobench/IsoBench) currently consists of 1,887 examples across four domains, though the construction of a larger dataset is currently underway. spanning diverse domains such as discrete and applied mathematics, physics, chemistry, and chess. For each sample, we evaluate multiple isomorphic input representations containing the same information—one with a visual representation and others with domain-specific textual representations—to facilitate multimodal performance assessments.
@@ 4 IsoCombination and IsoScratchPad
In many scientific fields, the images are complicated, researchers usually provide extra text inputs to help the model. This is also a common practice in scientific fields, where figures and tables in most papers are coupled with captions (for example, this paper), providing descriptions and takeaways. Motivated by these and based on the observations made above on IsoBench, we can design two simple and deliberately contrived mechanisms to further study the performance of multimodal foundation models on tasks with isomorphic representations: IsoCombination (IsoCB) and IsoScratchPad (IsoSP). IsoCB feeds multiple isomorphic representations to the multimodal model simultaneously in order to see the combined effect of using multiple representations on performance. Alternatively, IsoSP uses a scratchpad to explicitly translate from visual to text representations (and then uses this text representation to solve a task), taking advantage of the empirically higher-performing text representations for visual data. In particular, for IsoSP we employ a two-step prompting strategy: initially, the model receives a prompt featuring a visual representation, which it is tasked to translate into a textual format; then, the model is prompted with this generated text representation to predict the output. We illustrate both methods in Figure 3.
@@ 5 Related Works
There has been a plethora of recent developments in vision-language foundation models; they can be broadly categorized by their methods for representing visual modalities. A representative approach involves tokenizing visual inputs to be jointly trained with language inputs (Yu et al., 2023; Google, 2023; McKinzie et al., 2024; Chen et al., 2022; 2023; Team, 2024, inter alia). Another line of work processes continuous visual features by directly projecting them to the language embedding space via a learnable function (Liu et al., 2023b; a; 2024a; Bavishi et al., 2023). While the design choices of some API-access models (OpenAI, 2023b; Anthropic, 2024; Google, 2023; Reka, 2024) remain largely unknown, most performant models use early fusion, the practice of integrating visual and language features at the input level of an autoregressive model. At the core of these design choices is the hardness in representing visual features, which has been reported by several early studies (McKinzie et al., 2024) to be the key bottleneck towards performant vision-language foundation models. IsoBench aims to supplement prior works, which have mostly focused on analyzing modeling methods, with a data-centric view by measuring performance discrepancies incurred by multimodal representations.
@@
Many vision-language datasets have recently been curated to assess the capabilities of multimodal foundation models (Yue et al., 2023; Lu et al., 2024; Zhang et al., 2024a; Fan et al., 2024; Zhang et al., 2024b). They often consist of a large collection of problems with varying levels of difficulty and usage of the visual inputs. As a result, they are suitable for assessing VLMs as a whole, but cannot attribute these performances to capabilities of their LLM backbones and improvements from downstream vision-language training. While several contemporaneous works (Fan et al., 2024; Zhang et al., 2024a; b) introduce a language-only subset of their vision-language prompts, they focus primarily on assessing certain areas of foundation models, such as deductive reasoning and problems of beyond-polynomial complexity. IsoBench is the first of its kind that offers a holistic and fine-grained analysis of the performance gap induced by variations in input modalities. For instance, a contemporaneous work, MathVerse (Zhang et al., 2024a), studies redundant, but not necessarily equivalent input representations on math problems whilst IsoBench studies the impact of isomorphic representations on a broader set of tasks.
@@
Similar to what IsoBench observes on multimodal foundation models’ sensitivity to input modalities, uni-modal language models also show sensitivity to input perturbation. For example, LLMs are shown to be sensitive to subtle changes in zero-shot and few-show settings (Sclar et al., 2023; Chang & Jia, 2023). Meanwhile, the sensitivity of Transformers, the backbone of foundation models, is also broadly studied (Bombari & Mondelli, 2024; Bhattamishra et al., 2023; Vasudeva et al., 2024; Kong et al., 2024). To address the sensitivity issues, methods including prompt design (Yoo et al., 2021; Mishra et al., 2021; Le et al., 2023; Liu et al., 2023c) and prompting-based training strategies (Jain et al., 2023; Guo et al., 2023) have been employed. IsoBench is one of the pioneers to study model sensitivity to input representations and modalities.
@@
Recent work has shown that Transformers, the backbone of multimodal foundation models, are capable of implementing various algorithms after fine-tuning on particular tasks (Khalighinejad et al., 2023; Garg et al., 2022; von Oswald et al., 2022; Hanna et al., 2023; Fu et al., 2023; Lee et al., 2023; Zhou et al., 2024) or with proper prompting schemes (Wei et al., 2023; Yao et al., 2023; Liu et al., 2024b). However, the extent to which foundation models are capable of solving problems involving complex reasoning, algorithms, or scientific awareness remains unclear.
@@ A.1 Mathematics
Images for math problems are generated with 300 DPI in JPEG format.
@@ A.2 Algorithms
Images for algorithm problems are generated with 300 DPI in JPEG format.
@@ A.3 Games
Images for Chess are the same as Lichess, and are converted to JPEG formats.
@@ A.4 Sciences
Images for Chess are the same as ScienceQA, and are converted to JPEG formats.
@@ Appendix C Full IsoBench Results
![[Uncaptioned image]](extracted/5798893/sections/examples/science/figures/chemistry_solution_e1_v.png)
![[Uncaptioned image]](extracted/5798893/sections/examples/science/figures/chemistry_solution_e1_t.png)
![[Uncaptioned image]](extracted/5798893/sections/examples/science/figures/chemistry_solution_e2_v.png)
![[Uncaptioned image]](extracted/5798893/sections/examples/science/figures/chemistry_solution_e2_t.png)
![[Uncaptioned image]](extracted/5798893/sections/examples/science/figures/chemistry_solution_e3_v.png)
![[Uncaptioned image]](extracted/5798893/sections/examples/science/figures/chemistry_solution_e3_t.png)
![[Uncaptioned image]](extracted/5798893/sections/examples/science/figures/physics_solution_e1_v.png)
![[Uncaptioned image]](extracted/5798893/sections/examples/science/figures/physics_solution_e1_t.png)
![[Uncaptioned image]](extracted/5798893/sections/examples/science/figures/physics_solution_e2_v.png)
![[Uncaptioned image]](extracted/5798893/sections/examples/science/figures/physics_solution_e2_t.png)